INAF LBC+LUCIFER queue observing, 2011 June 27 -28

Observer: R. Speziali, K. Boutsia
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LUCIFER

Summary

In the afternoon there were clouds at the horizon.
Therefore we decided to start with a short program and high seeing limit (2"): the SN2011bn. Unfortunately there were two sources close to the galaxy and we decided to postpone the program after the verification with the PI.
Then we moved to Maiolino's field - but there are clouds passing by. The seeing is acceptable and we see two spectra in the slits as in the previous nights.
At midnight we moved to Magrini. At the the end of the 2nd hr in J, the preset failed because of clouds.
We have waited open, just in case the clouds would go away, sending the telluric preset. We did obtain a spectra for the telluric but with low counts.
We closed for clouds at around ~11:00UT - we consider the last 2hrs ( 9:45 UT that the preset failed up to 11:45 UT that is when we finished observing the previous nights) lost to clouds.

Data Summary

Project Objects Filter ExpTime Notes
Maiolino HDF 210_zJHK - H 1hr55min arc + flat + Tell
Magrini 7c17_cluster 210_zJHK - J 1hr45min arc+flat + Tell (counts are low because of clouds)
STD_tell HD121409 200_H+K - OrderSep 2min telluric for the SN2011bm which was not observed in the end

Details

02:48 - Correct pointing

03:12 - STD_Tell_SpecAcquisitionTemplate.txt 0001
  • preset failed because the guide star is slewing outside the GCS image
  • send preset again

03:14 - STD_Tell_SpecAcquisitionTemplate.txt
  • preset failed again

03:16 - STD_Tell_SpecAcquisitionTemplate.txt 0002-0004

03:26 - STD_Tell_SpecScienceTemplate.txt 0005-0006
  • took an image outside the scirpt just to make sure the telluric was still centered 0007

03:28 - SN2011bm_SpecAcquisitionTemplate.txt 0008-0010
  • we see 2 sources next to the galaxy
  • we are not sure which one is the SN - we need to contact the PI

03:47 - Slew to Tell_acq_HDF_H1 (Maiolino) 0011-0015

04:03 - Tell_int_HDF_H1 0016-0022

04:10 - Maiolino_Spec_Acquisition_HDF_H1 0023-0027
  • it is very difficult to center the mask because the sources are very faint
  • we can see only on source in a slit

04:34 - Maiolino_SpecScience_HDF_H1_3.txt 0028-0032
  • we are not sure that the mask is centered well so we interrupt the script after the 5th offset
  • we try MOS Acq again

05:02 - Maiolino_Spec_Acquisition_HDF_H1 - preset failed
  • pointing correction

05:04 - Maiolino_Spec_Acquisition_HDF_H1 0033-0036
  • the centering did not change much in comparison to the previous attempt
  • we decide to send the science scripts

05:29 - Maiolino_SpecScience_HDF_H1_3.txt 0037-0048
  • Y=-1.0 to center the mask

06:35 - Maiolino_SpecScience_HDF_H1_4.txt 0049-0054
  • the seeing is getting worse - there is also a fire nearby - there might be smoke in the air
  • anyway we can see an emission line in one slit and a continuum in another slit - so we keep going

07:08 - interrupt script after the 6th iteration because we see no spectra and the elevation has reached the PI limit
  • Calib Unit IN for calibrations
  • ARC 0055-0062
  • FLAT 0063-0068

07:21 - Calib Unit OUT

07:22 - magrini_acq_sci.txt 0069-0074
  • we manually re-adjust the mask after the MOS acquisition

07:48 - magrini_spec_sci_J_4.txt 0075-0086
  • DIMM seeing 0.8
  • there seems to be less smoke ...

08:54 - magrini_spec_sci_J_5.txt 0087-0095
  • during the 9th offset - we lost the guiding star - clouds, smoke ????

09:45 - Calib Unit IN
  • ARC 0096-0099
  • FLAT 0100-0114

09:57 - Calib Unit OUT - sending magrini_acq_PiHer.txt
  • preset failed - no star found because of clouds

10:00 - magrini_acq_PiHer.txt - preset failed again
  • we wait for the clouds to clear

10:19 - magrini_acq_PiHer.txt
  • preset failed - we don't know if it is clouds or pointing
  • pointing correction
  • it is probably clouds because we used a 3.5 mag star for pointing correction and it seemed faint

10:30 - magrini_acq_PiHer.txt - preset failed

10:48 - magrini_acq_PiHer.txt 0115-0117

10:56 - magrini_tell_int_PiHer_J.txt 0118-0121
  • actually we lost the the guide star but we turned off guiding and we too the images anyway
  • the counts are lowish (~1000count) - exptime 4sec
  • we repeated the exposure because because it seemed to clear - but counts are low again 0122-0125

11:07 - close dome for clouds
  • we consider the last 2 hours (9:45 UT that the preset failed up to 11:45 UT that is when we finished observing the previous nights) lost to clouds

-- KonstantinaBoutsia - 28 Jun 2011
Topic revision: r3 - 29 Jun 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback